# Exploring Cross-model Neuronal Correlations in the Context of Predicting Model Performance and Generalizability

### v2

@@ A. Small neural networks
We evaluated our method on five simple fully connected networks (FCNs) and a two-layer convolutional neural network (CNN). The number of neurons considered for each hidden layer is described in Table 1. These networks were trained on four datasets including MNIST (LeCun 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), Kuzushiji-MNIST (Clanuwat et al. 2018), and Afro-MNIST (Wu, Yang, and Prabhu 2020). All these datasets contain 60000 training examples and 10000 test examples. Each example comprises a $28\times 28$ image paired with a classification label from a set of 10 categories. MNIST, Fashion-MNIST, Kuzushiji-MNIST, and Afro-MNIST contain images of handwritten numerals, fashion, Hiragana characters, and African numbers respectively. The selection of datasets with similar attributes was deliberate as our focus was on understanding the interplay between networks, irrespective of data size or category.
@@
Table 3 illustrates the robustness of FCNs against the black-box attack detailed in Algorithm 1. Effect of the attack on some test images is illustrated in Figure 1. Results for CNN are also provided, showcasing its inferior robustness, given its role in generating the adversaries. In our specific scenario, we selected a norm value of $\epsilon=\frac{10}{255}$ and $N_{ep}=10$. The results indicate that, on average, Net5, Net2, Net4, Net3, and Net1 ranked from vulnerable to robust. Among the FCNs, Net5 exhibited the lowest overall accuracy. Notably, according to Table 2(d), Net5 displayed the weakest correlation with the other networks, aligning with its subpar performance. Conversely, Net3 demonstrated the highest average correlation score with the other networks, potentially attributed to its architectural similarity with the other FCNs.
Table 4: Correlation between CNN and other networks.
Dataset Net1 Net2 Net3 Net4 Net5
MNIST 0.880 0.856 0.908 0.831 0.416
Fashion-MNIST 0.756 0.822 0.820 0.755 0.317
Kuzushiji-MNIST 0.900 0.769 0.910 0.894 0.519
Afro-MNIST 0.588 0.629 0.730 0.718 0.059
@@ B. Large neural networks
The proposed correlation metric was also investigated on different variants of ResNet, including ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 (He et al. 2016). We used the pretrained weights on ImageNet (Deng et al. 2009) as our baselines. Since investigating the correlation on all of the neurons in all layers was not possible we only tested the method on the first output of the fourth layer on all ResNets. Also, we used only 10 test data points from the validation set of ImageNet. Considering the absolute values of the correlations, results suggest that the most similar networks to ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 are ResNet-34, ResNet-18, ResNet-101, ResNet-152, and ResNet-101. This pattern illustrates that networks with a similar number of layers exhibit stronger partial correlations, serving as a validation of the effectiveness of our proposed correlation metric.

### v3

@@ Methods
Table 3: Correlation between CNN and other networks.
Dataset Net1 Net2 Net3 Net4 Net5
MNIST 0.880 0.856 0.908 0.831 0.416
Fashion-MNIST 0.756 0.822 0.820 0.755 0.317
Kuzushiji-MNIST 0.900 0.769 0.910 0.894 0.519
Afro-MNIST 0.588 0.629 0.730 0.718 0.059
Numerical Results
@@ A. Small neural networks
We evaluated our method on five simple fully connected networks (FCNs) and a two-layer convolutional neural network (CNN). The number of neurons considered for each hidden layer is described in Table 1. These networks were trained on four datasets including MNIST (LeCun 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), Kuzushiji-MNIST (Clanuwat et al. 2018), and Afro-MNIST (Wu, Yang, and Prabhu 2020). All these datasets are comprised of 60000 training examples and 10000 test examples. Each example consists of a $28\times 28$ image paired with a classification label from a set of 10 categories. MNIST, Fashion-MNIST, Kuzushiji-MNIST, and Afro-MNIST contain images of handwritten numerals, fashion, Hiragana characters, and African numbers respectively. The selection of datasets with similar attributes was deliberate as our focus was on understanding the interplay between networks, irrespective of data size or category.
@@
Table 4 depicts the robustness of FCNs against the black-box attack detailed in Algorithm 1. The impact of the attack on some test images is illustrated in Figure 1. Results for CNN are also provided, showcasing its inferior robustness, given its role in generating the adversaries. In our specific scenario, we selected a norm value of $\epsilon=\frac{10}{255}$ and $N_{ep}=10$. The results indicate that, on average, Net5, Net2, Net4, Net3, and Net1 ranked from vulnerable to robust. Among the FCNs, Net5 exhibited the lowest overall accuracy. Notably, according to Table 2(d), Net5 displayed the weakest correlation with the other networks, aligning with its subpar performance. Conversely, Net3 demonstrated the highest average correlation score with the other networks, potentially attributed to its architectural similarity with the other FCNs.
@@ B. Large neural networks
The proposed correlation metric was also investigated on large neural networks such as ResNet (He et al. 2016) and DenseNet (Huang et al. 2017). ResNet, short for Residual Network, is a deep neural network architecture that leverages residual blocks and skip connections to enhance training efficiency and performance. The variants of ResNet include ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 (He et al. 2016). DenseNet, short for Dense Convolutional Network, is a deep neural network architecture that leverages dense connections between layers to promote feature reuse and efficient gradient flow, resulting in compact models that excel in various computer vision tasks. The architecture has several variants, namely DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201 (Huang et al. 2017). For both cases, we utilized the pretrained weights on ImageNet (Deng et al. 2009) as our baselines. We limited our analysis to just 10 test data points sampled from the validation set of ImageNet due to the time and space complexities inherent in our methodology. Since investigating the correlation on all of the neurons in all layers was not possible, we focused our analysis solely on the initial output of the fourth layer for all ResNets and the primary output of the third transition layer for all DenseNets. A point to be highlighted is that assessing the correlation metric on the final layers often yields superior accuracy compared to other layers, given that these layers encapsulate more profound and meaningful representations. Considering the absolute values of the correlations, the results shown in Table 5 suggest that the most similar networks to ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 are ResNet-34, ResNet-18, ResNet-101, ResNet-152, and ResNet-101 respectively. As demonstrated in Table 6, DenseNet-169, DenseNet-169, DenseNet-201, and DenseNet-169 have the most similarities to DenseNet-121, DenseNet-161, DenseNet-169, and DenseNet-201. This pattern illustrates that networks with roughly the same number of layers often exhibit more profound partial correlations, serving as a validation of the effectiveness of our proposed correlation metric. Furthermore, deeper neural networks exhibit greater robustness against adversarial attacks. Thus, our findings further reveal that a stronger correlation with a robust model correlates with higher robustness.
Table 6: Partial correlation between DenseNets.
DenseNet-121 DenseNet-161 DenseNet-169 DenseNet-201
DenseNet-121 1 0.657 0.664 0.637
DenseNet-161 X 1 0.688 0.631
DenseNet-169 X X 1 0.696
DenseNet-201 X X X 1
@@ Discussion
  • Layer depth impact: Evaluation of the partial correlation metric on the earlier layers is typically less accurate since the final layers tend to capture more sophisticated and significant representations of the data.